# Cognitive-Inspired Hierarchical Attention Fusion With Visual and Textual for Cross-Domain Sequential Recommendation

### v4

@@ Abstract
Introduction
Sequential Recommendation (SR) has emerged as a prominent approach for modeling dynamic user preferences by analyzing historical interaction sequences (?, ?, ?, ?, ?). While SR focuses on predicting the next item of interest within a single domain, it often suffers from data sparsity and domain bias, limiting its ability to capture the full spectrum of user behavior. To address these limitations, Cross-Domain Sequential Recommendation (CDSR) (?, ?, ?, ?, ?, ?, ?) has been proposed, leveraging information from multiple domains to enhance recommendation accuracy and provide a more comprehensive understanding of user preferences.
@@ Introduction
Early efforts in CDSR, such as $\pi$-Net (?, ?), model item interaction sequences within a single domain and transfer learned representations to other domains through gated transfer modules. Subsequent work, like MIFN (?, ?), incorporates external knowledge graphs to strengthen cross-domain connections. However, these approaches primarily focus on intra-sequence relationships within individual domains, often neglecting inter-sequence relationships across domains. This oversight leads to persistent domain bias and suboptimal performance.
Moreover, existing CDSR methods fail to fully exploit the rich visual information associated with items. Human decision-making is inherently multimodal, with visual cues playing a critical role in shaping user preferences (?, ?). For instance, users often form initial impressions of items based on their visual appearance, even when textual descriptions are similar. This observation aligns with cognitive theories emphasizing the integration of multimodal stimuli in preference formation (?, ?). To bridge this gap, we propose Hierarchical Attention Fusion of Visual and Textual Representations (HAF-VT), a novel framework that integrates visual and textual embeddings to enrich item representations and model cross-domain user preferences. By leveraging the frozen CLIP model (?, ?), we generate aligned image and text embeddings, capturing both intra- and inter-sequence relationships. Our hierarchical attention mechanism jointly learns single-domain and cross-domain preferences, mimicking the human ability to integrate information across contexts.
@@
Related Work
Models for Cognitive Science
Recent advancements in computer vision and machine learning (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?) have demonstrated significant potential for addressing key research questions in cognitive science (?, ?, ?, ?, ?, ?, ?, ?). A growing body of work has focused on image similarity and categorization (?, ?, ?, ?, ?, ?, ?), which, while related to the phenomenon of naming, does not directly address it. These studies collectively reveal that deep visual representations, particularly those derived from the final layers of neural networks, align closely with human perceptual judgments. This alignment is often measured by comparing human similarity ratings with model-generated similarity scores, highlighting the ability of deep learning models to capture human-like visual representations. More directly relevant to the present work are studies by (?, ?, ?, ?, ?), which leverage deep learning-based computer vision models to construct visual semantic representations for extensive vocabularies. These approaches bridge the gap between visual and linguistic domains, offering insights into how visual information can be mapped to semantic structures. Recent research has also begun to explore cognitive modeling using the CLIP (Contrastive Language–Image Pretraining) framework. The integration of self-supervised learning with vision-language pretraining has spurred the development of numerous CLIP-based models, each aiming to enhance the alignment between visual and textual data (?, ?). These models extend the original CLIP architecture by introducing novel techniques to improve representation quality and strengthen the connection between images and text.
@@
Sequential Recommendation
Sequential recommendation (?, ?) serves as the core framework of our approach, where user behavior is modeled as a sequence of time-sensitive items with the goal of predicting the likelihood of the next item in the sequence. Initially, Markov chains were utilized to capture the sequential dependencies within the item sequences, as seen in FPMC (?, ?). To capture more complex interactions between historical items, deep learning techniques such as recurrent neural networks (?, ?, ?), convolutional neural networks (?, ?), and attention mechanisms (?, ?) have been integrated into recommender systems (?, ?, ?, ?, ?, ?). In addition, SURGE (?, ?) uses metric learning to reduce the complexity of the item sequences. More recently, works like DFAR (?, ?) and DCN (?, ?) have explored capturing deeper, more intricate relationships within sequential recommendation data. In this work, we extend these ideas by applying cross-domain learning to sequential recommendation models, facilitating knowledge transfer between multiple domains.
Cross-Domain Recommendation
Cross-domain recommender systems (?, ?) have emerged as a promising approach to address the challenges of sparse data and cold-start issues typically faced in sequential recommendation tasks (?, ?). Early models in this area were grounded in single-domain recommendation frameworks, under the assumption that auxiliary user behaviors from different domains could contribute to improved user modeling in the target domain (?, ?). Many popular methods rely on transfer learning (?, ?), which involves transferring user or item embeddings from the source domain to the target domain to enhance the modeling process. Notable examples of such approaches include MiNet (?, ?), CoNet (?, ?), and itemCST (?, ?), among others. However, in practice, industrial platforms typically aim to improve all product domains simultaneously, rather than exclusively enhancing the target domain without considering the source domain. This has led to the development of dual learning (?, ?, ?), which enables mutual improvements across both the source and target domains, and has already been successfully applied in cross-domain recommender systems. Despite these advancements, many existing methods fail to account for the role of visual perceptual features and do not integrate multimodal fusion with attention mechanisms, which could significantly enhance performance.
Methodology
Problem Formulation
@@
Visual and Textual Feature Integration
@@
Hierarchical Attention in Preference Formation
In traditional sequential recommendation methods (?, ?, ?), behavioral sequences from domains $X$ and $Y$ are often processed indiscriminately, leading to potential dominance by the domain with higher behavioral frequency. This imbalance can skew the representation of user preferences. To address this issue, we propose a hierarchical attentional framework that separately processes sequences $S^{X}$, $S^{Y}$, and $S^{X+Y}$, ensuring balanced representation across domains. Our approach integrates item identity (ID), visual features, and textual embeddings to model user preferences more comprehensively. As illustrated in Fig. 3, to model the interactions within each sequence more effectively, we incorporate a self-attention mechanism with query, key, and value (QKV) pairs. The attention mechanism allows the model to focus on relevant parts of the sequence when making predictions, capturing both local and global dependencies within the sequence.
@@
$\mathrm{P}^{X}_{id}(x_{t+1}\mid\mathcal{S})=\text{softmax}\left(\text{Sim}(h_{id}^{X},E_{id}^{X})\right),\quad x_{t}\in S^{X},$ (4)
@@
$\mathrm{P}^{X}_{img}(x_{t+1}\mid\mathcal{S})=\text{softmax}\left(\text{Sim}(h_{img}^{X},E_{img}^{X})\right),\quad x_{t}\in S^{X},$ (5)
$\mathrm{P}^{X}_{tex}(x_{t+1}\mid\mathcal{S})=\text{softmax}\left(\text{Sim}(h_{tex}^{X},E_{tex}^{X})\right),\quad x_{t}\in S^{X},$ (6)
@@
$\mathcal{L}^{X}=\sum_{x_{t}\in S^{X}}-\log\mathrm{P}^{X}(x_{t+1}\mid\mathcal{S}).$ (8)
@@
$\mathrm{P}(x_{i}|S)=\mathcal{P}^{X}(x_{i}|S)+\lambda_{1}\mathcal{P}^{Y}(x_{i}|S)+\lambda_{2}\mathcal{P}^{X+Y}(x_{i}|S).$ (10)
@@
Experiments
In this section, we introduce the experimental setup, providing detailed descriptions of the dataset, evaluation metrics, and implementation specifics. We then compare our method with state-of-the-art approaches on the Amazon dataset (?, ?). Finally, we conduct ablation studies to validate the effectiveness of the proposed method.
Dataset and Evaluated Metric. Our experiments are conducted on the Amazon dataset (?, ?) to construct the CDSR scenarios. Following previous works (?, ?, ?), we select four domains to generate two CDSR scenarios for our experiments: “Food-Kitchen” and “Movie-Book”. We first extract users who have interactions in both domains. Then we filter out users and items with fewer than 10 interactions. Additionally, to meet the sequential constraints, we retain cross-domain interaction sequences that contain at least three items from each domain. In the training/validation/test split, the latest interaction sequences are equally divided into the validation and test sets, while the remaining interaction sequences are used for the training set. The statistics of our refined datasets for the CDSR scenarios are summarized in Table 1. We adopt Mean Reciprocal Rank (MRR) (?, ?) and Normalized Discounted Cumulative Gain (NDCG@){5, 10} (?, ?) as evaluation metrics to assess the performance of our model. These metrics are widely used in recommendation systems.
Table 1: Statistics of two CDSR scenarios.
@@
Implementation Details. For fair comparisons, we adopt the same hyperparameter settings as in previous works (?, ?). We set the learnable embedding size $q$ to 256, the CLIP-based image and text embedding size $e$ to 512, and the mini-batch size to 256, with a dropout rate of 0.3. The $L_{2}$ regularization coefficient is selected from $\{0.0001,0.00005,0.00001\}$, and the learning rate is chosen from $\{0.001,0.0005,0.0001\}$. Training is conducted for 100 epochs using an NVIDIA 4090 GPU, with the Adam optimizer (?, ?) to update all parameters.
Table 2: Experimental results (%) on the Food-Kitchen scenario.
@@
Table 3: Experimental results (%) on the Movie-Book scenario.
@@
GRU4Rec (?, ?) 3.83 3.14 3.73 1.68 1.34 1.52
SASRec (?, ?) 3.79 3.23 3.69 1.81 1.41 1.71
SR-GNN (?, ?) 3.85 3.27 3.78 1.78 1.40 1.66
PSJNet (?, ?) 4.63 4.06 4.76 2.44 2.07 2.35
MIFN (?, ?) 5.05 4.21 5.20 2.51 2.12 2.31
Tri-CDR (?, ?) 5.15 4.62 5.05 2.32 2.08 2.22
@@
Performance Comparisons
We evaluate our method against existing state-of-the-art (SOTA) approaches on the “Food-Kitchen” and “Movie-Book” CDSR scenarios, as summarized in Tab. 2 and Tab. 3, respectively. Our approach achieves superior performance, surpassing the current SOTA methods in terms of final evaluation metrics, thereby demonstrating its effectiveness in cross-domain recommendation tasks.
Table 4: Ablation study on the movie dataset
@@
Ablation Studies
We conducted an ablation study to systematically evaluate the contributions of our proposed components: Textual Fusion, Visual Fusion, and Hierarchical Attention. As illustrated in Tab. 4, we first established a baseline using the original framework without any additional modules, achieving an MRR of 5.03. The introduction of Visual Fusion alone resulted in a performance gain of 0.6 points, highlighting the importance of incorporating visual signals for cross-domain recommendation tasks. Similarly, adding Textual Fusion without hierarchical attention improved the MRR to 5.48, demonstrating the complementary role of textual information in enhancing the model’s understanding of user preferences.

### v5

@@ Abstract
I Introduction
Sequential Recommendation (SR) has emerged as a prominent approach for modeling dynamic user preferences by analyzing historical interaction sequences [1, 2, 3, 4]. While SR focuses on predicting the next item of interest within a single domain, it often suffers from data sparsity and domain bias, limiting its ability to capture the full spectrum of user behavior. To address these limitations, Cross-Domain Sequential Recommendation (CDSR) [5, 6, 7, 8, 9, 10] has been proposed, leveraging information from multiple domains to enhance recommendation accuracy and provide a more comprehensive understanding of user preferences.
@@ I Introduction
Early efforts in CDSR, such as $\pi$-Net [5], model item interaction sequences within a single domain and transfer learned representations to other domains through gated transfer modules. Subsequent work, like MIFN [11], incorporates external knowledge graphs to strengthen cross-domain connections. However, these approaches primarily focus on intra-sequence relationships within individual domains, often neglecting inter-sequence relationships across domains. This oversight leads to persistent domain bias and suboptimal performance.
Moreover, existing CDSR methods fail to fully exploit the rich visual information associated with items. Human decision-making is inherently multimodal, with visual cues playing a critical role in shaping user preferences [12]. For instance, users often form initial impressions of items based on their visual appearance, even when textual descriptions are similar. This observation aligns with cognitive theories emphasizing the integration of multimodal stimuli in preference formation [13]. To bridge this gap, we propose Hierarchical Attention Fusion of Visual and Textual Representations (HAF-VT), a novel framework that integrates visual and textual embeddings to enrich item representations and model cross-domain user preferences. By leveraging the frozen CLIP model [14], we generate aligned image and text embeddings, capturing both intra- and inter-sequence relationships. Our hierarchical attention mechanism jointly learns single-domain and cross-domain preferences, mimicking the human ability to integrate information across contexts.
@@
II Related Work
II-A Models for Cognitive Science
Recent advancements in computer vision and machine learning [15, 16, 17, 18, 19, 20, 21, 22, 23, 24] have demonstrated significant potential for addressing key research questions in cognitive science [25, 26, 27, 28, 29, 30, 31]. A growing body of work has focused on image similarity and categorization [32, 33, 34, 35, 36, 37], which, while related to the phenomenon of naming, does not directly address it. These studies collectively reveal that deep visual representations, particularly those derived from the final layers of neural networks, align closely with human perceptual judgments. This alignment is often measured by comparing human similarity ratings with model-generated similarity scores, highlighting the ability of deep learning models to capture human-like visual representations. More directly relevant to the present work are studies by [38, 39, 40, 41], which leverage deep learning-based computer vision models to construct visual semantic representations for extensive vocabularies. These approaches bridge the gap between visual and linguistic domains, offering insights into how visual information can be mapped to semantic structures. Recent research has also begun to explore cognitive modeling using the CLIP (Contrastive Language–Image Pretraining) framework. The integration of self-supervised learning with vision-language pretraining has spurred the development of numerous CLIP-based models, each aiming to enhance the alignment between visual and textual data [42]. These models extend the original CLIP architecture by introducing novel techniques to improve representation quality and strengthen the connection between images and text.
@@ II-A Models for Cognitive Science
II-B Sequential Recommendation
Sequential recommendation [43] serves as the core framework of our approach, where user behavior is modeled as a sequence of time-sensitive items with the goal of predicting the likelihood of the next item in the sequence. Initially, Markov chains were utilized to capture the sequential dependencies within the item sequences, as seen in FPMC [44]. To capture more complex interactions between historical items, deep learning techniques such as recurrent neural networks [45, 46], convolutional neural networks [47], and attention mechanisms [48] have been integrated into recommender systems [49, 50, 51, 52, 53]. In addition, SURGE [54] uses metric learning to reduce the complexity of the item sequences. More recently, works like DFAR [55] and DCN [56] have explored capturing deeper, more intricate relationships within sequential recommendation data. In this work, we extend these ideas by applying cross-domain learning to sequential recommendation models, facilitating knowledge transfer between multiple domains.
II-C Cross-Domain Recommendation
Cross-domain recommender systems [57] have emerged as a promising approach to address the challenges of sparse data and cold-start issues typically faced in sequential recommendation tasks [58, 59]. Early models in this area were grounded in single-domain recommendation frameworks, under the assumption that auxiliary user behaviors from different domains could contribute to improved user modeling in the target domain [60]. Many popular methods rely on transfer learning [61], which involves transferring user or item embeddings from the source domain to the target domain to enhance the modeling process. Notable examples of such approaches include MiNet [62], CoNet [63], and itemCST [64], among others. However, in practice, industrial platforms typically aim to improve all product domains simultaneously, rather than exclusively enhancing the target domain without considering the source domain. This has led to the development of dual learning [65, 66], which enables mutual improvements across both the source and target domains, and has already been successfully applied in cross-domain recommender systems. Despite these advancements, many existing methods fail to account for the role of visual perceptual features and do not integrate multimodal fusion with attention mechanisms, which could significantly enhance performance.
III Methodology
III-A Problem Formulation
@@ III-B Overall framework
III-C Visual and Textual Feature Integration
@@
III-D Hierarchical Attention in Preference Formation
In traditional sequential recommendation methods [49, 67], behavioral sequences from domains $X$ and $Y$ are often processed indiscriminately, leading to potential dominance by the domain with higher behavioral frequency. This imbalance can skew the representation of user preferences. To address this issue, we propose a hierarchical attentional framework that separately processes sequences $S^{X}$, $S^{Y}$, and $S^{X+Y}$, ensuring balanced representation across domains. Our approach integrates item identity (ID), visual features, and textual embeddings to model user preferences more comprehensively. As illustrated in Fig. 3, to model the interactions within each sequence more effectively, we incorporate a self-attention mechanism with query, key, and value (QKV) pairs. The attention mechanism allows the model to focus on relevant parts of the sequence when making predictions, capturing both local and global dependencies within the sequence.
@@ III-D Hierarchical Attention in Preference Formation
$\small\mathrm{P}^{X}_{id}(x_{t+1}\mid\mathcal{S})=\text{softmax}\left(\text{Sim}(h_{id}^{X},E_{id}^{X})\right),\quad x_{t}\in S^{X},$ (4)
@@
$\small\mathrm{P}^{X}_{img}(x_{t+1}\mid\mathcal{S})=\text{softmax}\left(\text{Sim}(h_{img}^{X},E_{img}^{X})\right),\quad x_{t}\in S^{X},$ (5)
$\small\mathrm{P}^{X}_{tex}(x_{t+1}\mid\mathcal{S})=\text{softmax}\left(\text{Sim}(h_{tex}^{X},E_{tex}^{X})\right),\quad x_{t}\in S^{X},$ (6)
@@
$\small\mathcal{L}^{X}=\sum_{x_{t}\in S^{X}}-\log\mathrm{P}^{X}(x_{t+1}\mid\mathcal{S}).$ (8)
@@
$\small\mathrm{P}(x_{i}|S)=\mathcal{P}^{X}(x_{i}|S)+\lambda_{1}\mathcal{P}^{Y}(x_{i}|S)+\lambda_{2}\mathcal{P}^{X+Y}(x_{i}|S).$ (10)
@@
IV Experiments
In this section, we introduce the experimental setup, providing detailed descriptions of the dataset, evaluation metrics, and implementation specifics. We then compare our method with state-of-the-art approaches on the Amazon dataset [68]. Finally, we conduct ablation studies to validate the effectiveness of the proposed method.
Dataset and Evaluated Metric. Our experiments are conducted on the Amazon dataset [68] to construct the CDSR scenarios. Following previous works [5, 11], we select four domains to generate two CDSR scenarios for our experiments: “Food-Kitchen” and “Movie-Book”. We first extract users who have interactions in both domains. Then we filter out users and items with fewer than 10 interactions. Additionally, to meet the sequential constraints, we retain cross-domain interaction sequences that contain at least three items from each domain. In the training/validation/test split, the latest interaction sequences are equally divided into the validation and test sets, while the remaining interaction sequences are used for the training set. The statistics of our refined datasets for the CDSR scenarios are summarized in Table I. We adopt Mean Reciprocal Rank (MRR) [69] and Normalized Discounted Cumulative Gain (NDCG@){5, 10} [70] as evaluation metrics to assess the performance of our model. These metrics are widely used in recommendation systems.
TABLE I: Statistics of two CDSR scenarios.
@@ IV Experiments
Implementation Details. For fair comparisons, we adopt the same hyperparameter settings as in previous works [11]. We set the learnable embedding size $q$ to 256, the CLIP-based image and text embedding size $e$ to 512, and the mini-batch size to 256, with a dropout rate of 0.3. The $L_{2}$ regularization coefficient is selected from $\{0.0001,0.00005,0.00001\}$, and the learning rate is chosen from $\{0.001,0.0005,0.0001\}$. Training is conducted for 100 epochs using an NVIDIA 4090 GPU, with the Adam optimizer [71] to update all parameters.
TABLE II: Experimental results (%) on the Food-Kitchen scenario.
@@
TABLE III: Experimental results (%) on the Movie-Book scenario.
@@
GRU4Rec [49] 3.83 3.14 3.73 1.68 1.34 1.52
SASRec [67] 3.79 3.23 3.69 1.81 1.41 1.71
SR-GNN [72] 3.85 3.27 3.78 1.78 1.40 1.66
PSJNet [74] 4.63 4.06 4.76 2.44 2.07 2.35
MIFN [11] 5.05 4.21 5.20 2.51 2.12 2.31
Tri-CDR [73] 5.15 4.62 5.05 2.32 2.08 2.22
@@
IV-A Performance Comparisons
We evaluate our method against existing state-of-the-art (SOTA) approaches on the “Food-Kitchen” and “Movie-Book” CDSR scenarios, as summarized in Tab. II and Tab. III, respectively. Our approach achieves superior performance, surpassing the current SOTA methods in terms of final evaluation metrics, thereby demonstrating its effectiveness in cross-domain recommendation tasks.
TABLE IV: Ablation study on the movie dataset
@@ IV-A Performance Comparisons
IV-B Ablation Studies
We conducted an ablation study to systematically evaluate the contributions of our proposed components: Textual Fusion, Visual Fusion, and Hierarchical Attention. As illustrated in Tab. IV, we first established a baseline using the original framework without any additional modules, achieving an MRR of 5.03. The introduction of Visual Fusion alone resulted in a performance gain of 0.6 points, highlighting the importance of incorporating visual signals for cross-domain recommendation tasks. Similarly, adding Textual Fusion without hierarchical attention improved the MRR to 5.48, demonstrating the complementary role of textual information in enhancing the model’s understanding of user preferences.